### v1

# Fractional Thouless pumping of solitons: a unique manifestation of bulk-edge correspondence of nonlinear eigenvalue problems

Abstract
Recent foundational studies have established the bulk-edge correspondence for nonlinear eigenvalue problems using auxiliary eigenvalues $\hat{H}\Psi=\omega S(\omega)\Psi$, spanning both linear [ T. Isobe et al., Phys. Rev. Lett. 132, 126601 (2024)] and nonlinear [ Chenxi Bai and Zhaoxin Liang, Phys. Rev. A. 111, 042201 (2025)] Hamiltionians. This raises a fundamental question: Can eigenvalue nonlinearity generate observable physical phenomena absent in conventional approaches ($\hat{H}\Psi=E\Psi$)? In this work, we demonstrate the first uniquely nonlinear manifestation of this correspondence: fractional Thouless pumping of solitons. Through systematic investigation of nonlinear Thouless pumping in an extended Rice-Mele model with next-nearest-neighbor (NNN) couplings, we discover that the NNN parameters can induce fractional topological phases despite quantized topological invariants from conventional approaches. Crucially, these fractional phases are explained within the auxiliary eigenvalue framework, linking nonlinear spectral features directly to bulk-boundary correspondence. This work reveals novel phenomena emerging from the interplay between nonlinearity and NNN couplings, providing key insights for designing topological insulators and controlling quantum edge states in nonlinear regimes.
preprint: APS/123-QED
I Introduction
The bulk-edge correspondence stands as a cornerstone of topological matter [1, 2, 3, 4, 5], establishing an intimate link between a system’s bulk topology and the emergence of robust boundary states [6, 7, 8, 9, 10] in diverse systems such as photonic systems [11, 12, 13, 14, 15, 16], classical wave systems [17, 18, 19, 20, 21, 22, 23], acoustic metamaterials [24], and even biological networks [25]. Central to this framework are topological invariants [26, 27, 28, 29, 30, 31]—Chern numbers, $\mathbb{Z}_{2}$ indices, and their generalizations—which serve as fundamental markers of bulk topology. These invariants are computed via integration over the Brillouin zone using methods like Wilson loops [32] or K-theory [33] or real-space wavefunction analysis, derived from the Schrödinger equation $\hat{H}\Psi=E\Psi$ and its linear eigenvalue solutions. In non-Hermitian systems [34, 35, 36, 37, 38, 39, 40], this paradigm undergoes a profound shift: the standard Brillouin zone must be replaced by a generalized Brillouin zone (GBZ) [41, 42, 43], a complex momentum contour, to reconcile bulk topology with skin effects [44, 45, 46] and predict edge states. The universality of bulk-edge correspondence finds its deepest expression in topological index theorems (e.g., Atiyah–Singer [47, 48, 49, 50, 51, 52, 53]), manifested through physical boundary phenomena. Future work can extend this foundation to unify bulk-edge physics across interacting [54, 55, 56, 57, 58, 59] disordered, and driven non-equilibrium systems [60, 61, 62]—a grand challenge demanding new mathematical tools.
Conceptually going beyond above conventional approaches, a recent breakthrough [63] has established a bridge between bulk-edge correspondence and eigenvalue nonlinearity in linear Hamiltonians throuth auxiliary eigenvalues $\hat{H}\Psi=\omega S(\omega)\Psi$. This framework demonstrates how topological edge states of auxiliary eigenstates emerge as physical edge states under weak but finite nonlinearity (requiring monotonic $\omega$-dependence), rigorously extending bulk-edge correspondence to nonlinear eigenvalue constraints [64, 65, 66, 67, 68, 69, 70, 71, 72]. Building on this foundation, our work [73] generalizes the auxiliary eigenvalue approach to intrinsic nonlinear Hamiltonians, revealing an anomalous bulk-edge correspondence rooted in these nonlinear eigenvalues. Collectively, Refs. [63, 73] provide a unified description of bulk-edge physics across linear and nonlinear systems via nonlinear eigenvalue problems. This raises a key question: Can this framework predict observable phenomena inaccessible to conventional linear bulk-edge paradigms?
This work explores bulk-edge correspondence under eigenvalue nonlinearity $\hat{H}_{\text{ERM}}\Psi=\omega S(\omega)\Psi$, yielding physical observables inaccessible to conventional linear correspondence ($\hat{H}_{\text{ERM}}\Psi=E\Psi$). We implement this through nonlinear topological phase diagrams in an extended Rice-Mele model with next-nearest-neighbor couplings (NNN) $\{t_{a},t_{b}\}$ and nonlinearity strength $g$. First, we present topological phase diagrams for both linear ($H_{\text{ERM}}\Psi=E\Psi$) and nonlinear ($\hat{H}_{\text{ERM}}\Psi=\omega S(\omega)\Psi$) regimes, showing Chern number $\mathcal{C}$ evolution (Figs. 1(b),(c)). Crucially, we identify fractional Chern numbers in specific parameter regimes (Fig. 1(c)). We then investigate soliton pumping dynamics by tracking center-of-mass evolution through adiabatic cycles by the numerical methods of dynamically evolved soltions and instantaneous solitons with the theoretical framework of nonlinear eigenvalue problem (see Figs. 2-4 ) and anomolous eigenvalue’s nonlinearity problem (see Figs. 5-7) respectively. Both approaches demonstrate how dynamically tuned NNN hoppings and nonlinearity $g$ drive integer-to-fractional Chern phase transitions (Figs. 4(e), 7(e)). Finally, by constructing a 2D momentum-cyclic parameter space, we define topological invariants correlated with nonlinear eigenvalue bifurcations, obtaining a Haldane-model-like phase diagram. Our results predict directly observable fractional nonlinear Thouless pumping in photonic waveguides or cold-atom systems.
Figure 1: (a) Schematic diagram of the nonlinear extended RM model. The dotted box indicates the unit cell composed of sublattices A and B. (b) Chern number $\mathcal{C}$ topological phase diagram of the linear extended RM Hamiltonian (1) with $g=0$ under $H_{\text{ERM}}\Psi=E\Psi$. (c) Chern number $\mathcal{C}$ topological phase diagram of the linear extended RM Hamiltonian (1) with $g=0$ under $H_{\text{ERM}}\Psi=\omega S(\omega)\Psi$. The red dot represent the fractional Chern number of $\mathcal{C}=-\frac{1}{2}$ in specific NNN coupling parameter regimes of $t_{a}$ and $t_{b}$. (b,c) The parameters are given as $J=1$, $\delta=0.5$, $\Delta=1$, $T=2000\pi$, and $\omega=10^{-3}$.
The organizational structure of this paper is outlined as follows: In Sec. II, we provide an in-depth introduction to the Hamiltonian and Schrödinger equation of the nonlinear extended RM model. In Sec. III, we investigate the energy spectra and phase diagrams of the nonlinear extended RM model under both linear and nonlinear eigenvalue conditions, examining the effects of nonlinearity strength and NNN parameters on these properties. Sec. IV discusses bulk-edge correspondence in the nonlinear eigenvalue regime, elucidating the role of model parameters. Finally, in Sec. V, we summarize the results of our investigation.
II Nonlinear extended Rice-Mele model
The model system studied in this work consists of a one-dimensional interacting bosonic chain with $N$ dimer units. As illustrated in Fig. 1 (a), each unit cell contains two lattice sites (denoted as A and B respectively). As shown in Fig. 1(a), each unit cell comprises two lattice sites (A and B). At the mean-field level, our model system can be well described by the nonlinear extended RM Hamiltonian [74, 75, 76, 77, 78, 79, 80, 56, 81, 82, 78, 83]
$H_{\text{ERM}}=H_{\text{RM}}+H_{\text{NNN}},$ (1)
with
[…]
-- $\Delta\cos\left(\omega t\right)\sum_{n=1}^{N}\left(|\Psi_{n,A}|^{2}-|\Psi_{n,B}|^{2}\right)$
-- $\frac{g}{2}\sum_{n=1}^{N}\left(\left|\Psi_{n,A}\right|^{4}+\left|\Psi_{n,B}\right|^{4}\right),$
[…]
In Eq. (2), $\Psi_{n}=[\psi_{n,A},\psi_{n,B}]$ denotes the wave function for the dimer at site $n$ (two atoms per unit cell). The intracell and intercell couplings are $J_{1}=-J-\delta\sin(\omega t)$ and $J_{2}=-J+\delta\sin(\omega t)$ respectively, where $J$ represents the uniform hopping amplitude and $\delta\sin(\omega t)$ describes the time-dependent dimerization order. The staggered sublattice potential $\Delta\cos(\omega t)$ introduces an energy offset between sites, while $g$ quantifies the strength of the focusing Kerr-like nonlinearity [76]. The NNN Hamiltonian $H_{\text{NNN}}$ contains hopping amplitudes $t_{a}$ and $t_{b}$ along sublattices A and B. Crucially, the model reduces to the standard nonlinear Rice-Mele Hamiltonian [76] when $t_{a}=t_{b}=0$. Finally, we note that Hamiltonian (2) provides a direct theoretical framework for recent experiments [56] investigating interaction-induced Thouless pumping in dynamical superlattices or pulsed light propagation in waveguide arrays [76].
II.1 Topological phase diagram for linear and nonlinear eigenvalue problems
The topological physics of the nonlinear extended Rice-Mele model [Hamiltonian (1)] is governed by five parameters: nearest-neighbor couplings $J_{1}$, $J_{2}$, NNN hoppings $t_{a}$, $t_{b}$, and interaction strength $g$. The emphasis and value of this work is to demonstrate that eigenvalue nonlinearity [Refs. [63, 73]] enables novel bulk-edge correspondence, yielding physical observables inaccessible in linear systems ($H_{\text{ERM}}\Psi=E\Psi$).
[…]
Next, there are two kinds way of defining the eigenvalue problems of Hamiltonian (1) [63, 73], i. e. one is based on the conventional approaches of $H_{\text{ERM}}\Psi=E\Psi$; the other is referred as to the the bulk-edge correspondence of nonlinear eigenvalue problems of linear Hamiltonian (1) with $g=0$ defined as [63, 73]
$H_{\text{ERM}}(k,t)\Psi=\omega S(\omega,k)\Psi.$ (4)
Then, we introduce the matrix $P\left(\omega,k\right)$ and $P\left(\omega,k\right)=H_{\text{ERM}}\left(\omega,k\right)-\omega S\left(\omega,k\right)$. The solution satisfying the equation $P\left(\omega,\mathbf{k}\right)\varPsi=0$ is equivalent to that of the nonlinear equation presented in Eq. (4). In order to gain a more profound understanding of the bulk-edge correspondence model encapsulated by this nonlinear equation, we introduce an auxiliary eigenvalue $\lambda$, where $\lambda$ is an element of the real numbers $\mathbb{R}$, and the equation is transformed into $P\left(\omega,k\right)\varPsi=\lambda\varPsi$. We remark that the auxiliary eigenvalue $\lambda$ does not carry physical significance in the general case, with the notable exception when $\lambda=0$. Hence, the core problem reduces to determining the solution of $P\left(\omega,k\right)\varPsi=\lambda\varPsi$ specifically at $\lambda=0$. At last, the overlap matrix $S$ appearing in Eq. (4), which exhibits a dependence on the nonlinear parameter $\omega$, is constructed as $S=I-M_{S}\sigma_{z}$ with $M_{S}\left(\omega\right)=M_{1}\tanh\left(\omega t\right)/\omega$ and $\sigma_{z}$ being the $z$-component of Pauli matrix.
Finally, in Figs. 1(b) and (c), we present the topological phase diagrams depicting the Chern number $\mathcal{C}$ for the linear eigenvalue problem ($H_{\text{ERM}}\Psi=E\Psi$) and the nonlinear eigenvalue problem (Eq. 4), respectively, both evaluated at $g=0$ for $H_{\text{ERM}}$. Within the linear framework ($H_{\text{ERM}}\Psi=E\Psi$), the Chern number $\mathcal{C}$ remains quantized at $1$ (Fig. 1(b)). In contrast, the nonlinear eigenvalue problem $H_{\text{ERM}}\Psi=\omega S(\omega)\Psi$ yields fractional $\mathcal{C}$, exemplified by $\mathcal{C}=-\frac{1}{2}$ at $t_{a}=0.5$, $t_{b}=-0.5$ (red dot, Fig. 1(c)). Our subsequent goal is to demonstrate how the bulk-edge correspondence, modified by the eigenvalue nonlinearity evident in Fig. 1(c), gives rise to the observable physical phenomenon of fractional soliton pumping. This phenomenon is absent in the linear case shown in Fig. 1(b).
II.2 Nonlinear extended Schrödinger equation
Our strategy for demonstrating the physical manifestations of the fractional Chern number within the nonlinear framework (Fig. 1(c)) is inspired by the concept of nonlinear Thouless pumping [55, 55, 84, 57, 59]. Specifically, for the nonlinear extended RM Hamiltonian (1) with $g\neq 0$, the nonlinearity induces quantization of transport via soliton formation and non-uniform band occupation. Consequently, solitons are expected to undergo fractional pumping in the scenario depicted in Fig. 1(c).
In order to investigate the dynamics of soliton, we proceed to obtain equations of motion of nonlinear extended RM Hamiltonian $H_{\text{ERM}}$ by variation of the Hamiltonian $H_{\text{ERM}}$ as $i\partial\Psi_{j}/\partial t=\delta H/\delta\Psi^{*}_{j}$ which are described by the following set of nonlinear Schrödinger equations for $j=1,...N-1,N$,
$i\frac{\partial\Psi_{j,A}}{\partial t}=-\left(J+\delta\sin\omega t\right)\Psi_{j,B}-\left(J-\delta\sin\omega t\right)\Psi_{j-1,B}$
$-\left[\Delta\cos\omega t+g\left|\Psi_{j,A}\right|^{2}\right]\Psi_{j,A}\!+\!t_{a}(\Psi_{j-1,A}\!+\!\Psi_{j+1,A}),$ (5)
$i\frac{\partial\Psi_{j,B}}{\partial t}=-\left(J+\delta\sin\omega t\right)\Psi_{j,A}-\left(J-\delta\sin\omega t\right)\Psi_{j-1,A}$
$+\left[\Delta\cos\omega t-g\left|\Psi_{j,B}\right|^{2}\right]\Psi_{j,B}\!+\!t_{b}(\Psi_{j-1,B}\!+\!\Psi_{j+1,B}).$ (6)
The nonlinear eigenvalue $\mu$ of $\hat{H}_{\text{ERM}}$, defined through $\hat{H}_{\text{ERM}}\Psi=\mu\Psi$, can be derived by substituting $\Psi\rightarrow\Psi e^{i\mu t/\hbar}$ into Eqs. (5) and (6). We emphasize that Eqs. (5) and (6) represent a form of time-dependent Gross-Pitaevskii equations—equivalently, discrete nonlinear Schrödinger equations. This framework powerfully describes both the dynamics of Bose-Einstein condensates in ultracold quantum gases [56] and pulsed light propagation in waveguide arrays [76].
Figure 2: (a1)-(d1): Band structures of the nonlinear eigenvalues for the nonlinear extended RM Hamiltonian (1), (a2)-(d2): The position expectation value $\langle X\rangle$ of the soliton as a function of time over a single period. The parameters are fixed as $J=1$, $\delta=0.5$, $\Delta=1$, $t_{a}=0.4$, $t_{b}=-0.1$, $T=2000\pi$, and $\omega=10^{-3}$. (a1)-(d1) [(a2)-(d2)]: The interaction strengths are set to $g=0$, $g=1$, $g=5$ and $g=10$, respectively.
Figure 3: NNN tunneling control of topological transport: (a,c) Nonlinear eigenvalue spectra ($\mu$); (b,d,e) Quantized soliton dynamics. Fixed parameters: $J=1$, $\delta=0.5$, $\Delta=1$, $g=1$, $T=2000\pi$, $\omega=10^{-3}$. Tunneling configurations: (a,b) $t_{a}=0.4$, $t_{b}=-0.1$; (c,d) $t_{a}=0.5$, $t_{b}=-0.5$.
III Fractional nonlinear Thouless pumping of nonlinear extended RM Hamiltonian
In Sec. II, we presented the nonlinear extended RM Hamiltonian (1) along with the associated discrete nonlinear Schrödinger equations (5) and (6). In Sec. III, we plan to investigate the quantized and fractionally-quantized nonlinear Thouless pumping based on Eqs. (5) and (6). In more details, The objective of subsection III is to present the band structures of the nonlinear extended RM model as well as the pumping behavior of the soliton, while investigating the effects of the parameters $t_{a}$ and $t_{b}$ in the NNN Hamiltonian term (3). In the subsequent subsection (III.2), we adopt the methodology detailed in Ref. [63] to delve into the anomalous nonlinearity of the eigenvalues within the nonlinear extended RM Hamiltonian (1) and its corresponding bulk-edge correspondence.
III.1 Nonlinear eigenvalue of the nonlinear Hamiltonian
In this subsection III.1, we are interested into both the quantized nonlinear Thouless pumping and fractional nonlinear Thouless pumping predicted by the topological phase diagram in Figs. 1(b) and (c). This is achieved by numerically solving Eqs. (5) and (6) under a continuous change of the nonlinearity parameter $g$ in Eq. (1) from zero to a finite value. All simulations employ a $100$-unit cell lattice ($200$ sites). The initial soliton state, localized at site $n=0$, is prepared using the iterative self-consistent algorithm from Refs. [73, 76] with trial wavefunction $\Psi_{0}=\text{sech}\left[\lvert x-100\rvert/5\right]$. Here, we use two distinct methodologies to solve Eqs. (5) and (6) numerically. In the first approach entitled the method of dynamically evolved soltions, we use a fourth-order Runge-Kutta algorithm to numerically propagate the initial soliton profile through time evolution governed by Eqs. (5) and (6). This method directly simulates the dynamical evolution of the wavefunction under physical time-dependent driving. In the second approach called by the method of instantaneous solitons, an iterative self-consistent scheme is employed to converge to the steady-state solution at each time slice $t$, bypassing explicit time propagation. Under adiabatic conditions ($\omega\rightarrow 0$), both methods yield identical solutions, providing mutual validation of their accuracy.
Now, we are ready to investigate soliton pumping dynamics by monitoring the center-of-mass position through one adiabatic cycle using the two aforementioned methods. This requires calculating the expectation value $\langle X\rangle=\sum_{j}j|\Psi_{j}|^{2}$.
First, we consider how a soliton can be pumped under periodic driving through the lens of nonlinear eigenvalues (see Figs. 2(b1)-(d1)) corresponding to $H_{\text{ERM}}\Psi=\mu\Psi$ in Table I in Ref. [73] with the fixed $t_{a}$ and $t_{b}$. For baseline comparison, we consider the noninteracting case ($g=0$) of Hamiltonian (1), whose band structure is shown in Fig. 2(a1). Here, solitons are naturally absent due to vanishing nonlinearity, and consequently, no soliton pumping occurs - as expected and demonstrated in Fig. 2(a2). With the introduction of nonlinearity ($g\neq 0$) in Hamiltonian (1), solitons emerge through spontaneous symmetry-breaking bifurcations [54] and subsequently manifest pumping behavior, as clearly demonstrated in Figs. 2(b2)-(d2).
The quantized soliton transport in Fig. 2 originates from the Chern number (see Fig. 1(b)) in the noninteracting limit ($g=0$) [54, 73, 76, 84]. Specifically, under weak nonlinearity (Fig. 2(b1)), the soliton occupies the lowest band of Hamiltonian (1), whose linear dynamical Chern number equals $1$ (Fig. 1(b)). Consequently, the bulk-edge correspondence for the nonlinear extended RM Hamiltonian (1) yields quantized displacement of two lattice sites per cycle (Fig. 2(b2)). In contrast, strong nonlinearity (Fig. 2(d2)) induces Rabi oscillations between the two lowest bands. The vanishing sum of dynamical Chern numbers ($\mathcal{C}_{1}+\mathcal{C}_{2}=0$) then freezes soliton transport. For moderate interactions $g$ (Fig. 2(c2)), transport remains quantized when calculated via the instantaneous soliton method. However, dynamically evolved soliton trajectories exhibit position jumps (Fig. 2(c2)) due to emergent self-intersecting bands in the intermediate nonlinear regime. This band topology, characterized by loop structures [85], is visualized in Fig. 2(c1).
We next examine how NNN tunneling amplitudes $t_{a}$ and $t_{b}$ in Hamiltonian (1) influence quantized soliton transport, maintaining fixed weak nonlinearity ($g=1$). These effects are systematically investigated through Fig. 3. We first analyze the case $t_{a}=0.4$, $t_{b}=-0.1$. As demonstrated in Fig. 3(a), gapless edge states persist due to preserved inversion symmetry, resulting in stable soliton oscillations (Fig. 3(b)). Crucially, the ground-state soliton exhibits a pumping value of $2$ at this parameter point—consistent with Chern number $\mathcal{C}=1$ (corresponding to two atoms per unit cell)—and displays perfect adiabatic following, where dynamically evolved solitons precisely match their instantaneous counterparts. When the absolute values of $t_{b}$ becomes comparable to $t_{a}$ ($t_{a}=0.5$, $t_{b}=-0.5$), inversion symmetry breaks. This symmetry breaking opens a bandgap in the $\mu$-spectrum (Fig. 3(c)) and hybridizes edge states into bulk modes, thereby disrupting the adiabatic pathway as evidenced in Fig. 3(d). By selecting distinct NNN tunneling amplitudes $\{t_{a},t_{b}\}$ and superposing their corresponding soliton trajectories as shown in Fig. 3(e), we observe that all pumping pathways exhibit perfect adiabatic following except along the symmetry-broken line $t_{a}=-t_{b}=0.5$. Crucially, adiabatic breakdown occurs exclusively on this $t_{a}=-t_{b}$ trajectory. This reveals how NNN tunneling parameters reconfigure the system’s topological phase diagram.
Figure 4: NNN tunneling modulation of topological phases: (a,c) Nonlinear eigenvalue spectra ($\mu$); (b,d,e) Quantized soliton transport dynamics. Fixed parameters: $J=1$, $\delta=0.5$, $\Delta=1$, $g=6.215$, $T=2000\pi$, $\omega=10^{-3}$. Parameter configurations: (a,b) $t_{a}=0.4$, $t_{b}=-0.1$; (c,d) $t_{a}=0.5$, $t_{b}=-0.5$.
Finally, we demonstrate how the bulk-edge correspondence—modified by nonlinear eigenvalue topology shown in Fig. 1(c)—gives rise to fractional soliton pumping. This phenomenon, absent in the linear regime (Fig. 1(b)), is investigated at $g=6.215$. As depicted in Fig. 4(d), the $t_{a}=-t_{b}=0.5$ configuration maintains adiabatic breakdown (first observed at $g=1$), corresponding to the ground-state soliton’s self-intersecting structure in Fig. 4(c). Crucially, the transport trajectory obtained via the instantaneous soliton method (blue dotted curve, Fig. 4(d)) reveals fractional quantized pumping of the soliton. Conversely, for $t_{a}=0.4$, $t_{b}=-0.1$, self-trapping emerges (Fig. 4(a)), suppressing soliton dynamics and consequently halting pumping (Fig. 4(b)).
Motivated by the signal of observed fractional Thouless pumping phenomena induced by NNN tunneling $\{t_{a},t_{b}\}$ in Fig. 4(d), we are motivated to plot Fig. 4(e), which directly demonstrates how the bulk-edge correspondence—modified through nonlinear eigenvalue topology in Fig. 1(c)—generates fractional soliton pumping. This constitutes the central finding of our work. The parametric dependence in Fig. 4(e) reveals distinct soliton transport regimes: three trajectories exhibit adiabatic breakdown, while configurations $t_{a}=t_{b}=0.3$ and $t_{a}=0.4$, $t_{b}=-0.1$ show dynamical suppression and consequent pumping cessation. Crucially, we observe the crystallization of fractional topological phases in the $t_{a}=-t_{b}=0.2$ and $t_{a}=-t_{b}=0.5$ trajectories. This demonstrates selective reconfiguration of the phase diagram through NNN tunneling, particularly along the $t_{a}=-t_{b}$ symmetry line. Systematic comparison of $t_{a}=-t_{b}=0.5$ configurations reveals nonlinear control of quantization: at $g=1$ the pumping value remains $2$, while increasing to $g=6.215$ reduces it precisely to $1$.
Figure 5: Eigenvalue’s nonlinearity of nonlinear extended RM Hamiltonian and nonlinear Thouless pumping of soliton. (a1)-(d1): The auxiliary $\lambda$-spectrum for the nonlinear extended RM Hamiltonian (1). (a2)-(d2): The anticipated position of the nonlinear excitation of soliton as a function of time over one period. The parameters are fixed at $J=1$, $\delta=0.5$, $\Delta=1$, $t_{a}=0.4$, $t_{b}=-0.1$, $T=2000\pi$, and $\omega=10^{-3}$. Specifically, (a1)-(d1) and their corresponding (a2)-(d2) panels represent interaction strengths of $g=0$, $g=1$, $g=5$ and $g=10$, respectively.
Figure 6: Effects of parameter $t_{a}$ and $t_{b}$ on the auxiliary $\lambda$-spectrum defined in Eq. (14) for the nonlinear extended RM Hamiltonian (1).The parameters are fixed at $J=1$, $\delta=0.5$, $\Delta=1$, $g=1$, $T=2000\pi$, and $\omega=10^{-3}$. Other parameters are given as (a-b): $t_{a}=0.4$, $t_{b}=-0.1$; (c-d): $t_{a}=0.5$, $t_{b}=-0.5$. (e): Effects of parameter $t_{a}$ and $t_{b}$ on the anticipated position of the nonlinear excitation of soliton as a function of time over one period.
III.2 Anomalous eigenvalue’s nonlinearity of nonlinear Hamiltonian
In the preceding subsection (III.1), we have analyzed both quantized and fractional nonlinear Thouless pumping for the extended RM Hamiltonian (1) within the linear eigenvalue framework. In the subsequent subsection (III.2), we employ the methodology of Refs. [63, 73] to explore anomalous spectral nonlinearities in Hamiltonian (1) and their implications for anomalous bulk-edge correspondence first pointed by Ref. [73].
Anomalous eigenvalue nonlinearity generalizes the concept from linear Hamiltonians [63] to nonlinear systems, defined through $\hat{H}_{\text{non}}\Psi=\omega S(\omega)\Psi$. This framework exhibits dual nonlinearity—both Hamiltonian $\hat{H}_{\text{non}}$ and eigenvalue structure $\omega S(\omega)$ become state-dependent—distinguishing it from systems with only wavefunction nonlinearity (e.g., $g|\Psi_{n}|^{2}$). Such dual nonlinearity captures quantum systems with anharmonic potentials where energy levels and wavefunctions nonlinearly intertwine, providing a unified description of strongly correlated quantum states.
Our strategy for examining this anomalous eigenvalue nonlinearity in the nonlinear extended RM Hamiltonian (1) and bulk-edge correspondence, given the underlying nonlinearity of the eigenvalues, involves the utilization of auxiliary eigenvalues, as outlined below:
(i) The anomalous eigenvalue’s nonlinearity of nonlinear extended RM Hamiltonian (1) is defined as [73]
$H_{\text{ERM}}\left(\omega,t\right)\varPsi=\omega S\left(\omega,t\right)\varPsi.$ (7)
Note that Eq. (7) together with $S$ are employed to introduce auxiliary eigenvalues and furthermore the topological edge states of auxiliary eigenstates are topologically inherited as physical edge states, which represents interplay between the topology and nonlinearity of eigenvalues. The overlap matrix $S$ appearing in Eq. (7), which exhibits a dependence on the nonlinear parameter $\omega$, is constructed in the following manner:
$S\left(\omega\right)=\left(\begin{array}[]{cccc}S_{0}&0&0&0\\ 0&S_{0}&0&0\\ 0&0&\ddots&0\\ 0&0&0&S_{0}\end{array}\right),$ (12)
with $S_{0}$ being a diagonal matrix $S_{0}=1-\sigma_{z}M_{S}$ and $\sigma_{z}$ the $z$-component of Pauli matrix. Here, $M_{S}\left(\omega\right)$ is defined as $M_{S}\left(\omega\right)=M_{1}\tanh\left(\omega t\right)/\omega$.
(ii) Next, we proceed to introduce the matrix $P\left(\omega,t\right)$ as follows:
$P\left(\omega,t\right)=H_{\text{ERM}}\left(\omega,t\right)-\omega S\left(\omega,t\right).$ (13)
From Equation (13), the solution satisfying the equation $P\left(\omega,\mathbf{k}\right)\varPsi=0$ is equivalent to that of the nonlinear equation presented in Eq. (7). In order to gain a more profound understanding of the bulk-edge correspondence model encapsulated by this nonlinear equation, we introduce an auxiliary eigenvalue $\lambda$, where $\lambda$ is an element of the real numbers $\mathbb{R}$,
$P\left(\omega,t\right)\varPsi=\lambda\varPsi.$ (14)
We remark that the auxiliary eigenvalue $\lambda$ does not carry physical significance in the general case, with the notable exception when $\lambda=0$. Hence, the core problem reduces to determining the solution of Eq. (14) specifically at $\lambda=0$. Before proceeding, it is worth noting two key distinctions in solving Eqs. (4) and (7). First, Eq. (4) describes the eigenvalue problem for the linear Hamiltonian 1 with $g=0$, whereas Eq. (7) corresponds to the nonlinear case with $g\neq 0$. Second, these equations are solved in different representations: Eq. (4) in momentum space and Eq. (7) in coordinate space.
(iii) Finally, we turn to investigating both quantized and fractional nonlinear Thouless pumping within the framework of anomalous bulk-edge correspondence [73] for the extended Rice-Mele Hamiltonian (1). This analysis will proceed by numerically solving the eigenvalue equation (14). Furthermore, we will demonstrate topological inheritance from auxiliary to physical systems, showing how topological edge states of the auxiliary eigenstates manifest as physical edge states through numerical solution of Eqs. (5) and (6). To systematically explore these phenomena, we examine three distinct scenarios.
In the first scenario, we examine the connection between the auxiliary eigenvalue $\lambda$ (Figs. 5(a1)-(d1)) obtained from Eq. (14) and the soliton transport dynamics (Figs. 5(a2)-(d2)) by fixing the $t_{a}=0.4$ and $t_{b}=-0.1$. Specifically, at $g=0$, the pump transport oscillates near the initial point, as shown in Fig. 5(a2). As the interaction strength increases to $g=1$ and $g=5$, the system demonstrates pump transport phenomena. When $g=1$, the pump transport value over one period remains 2, as depicted in Fig. 5(b2). Furthermore, when the interaction strength continues to rise to $g=5$, the dynamical evolution of the soliton wave within a period is significantly reduced, and the pumping value decreases from 2 to 1, as depicted in Fig. 5(c2). Consistent with the original nonlinear extended RM model, when $g=10$, the pumping ceases to operate in the nonlinear extended RM model with the nonlinear eigenvalue problem, as illustrated in Fig. 5(d2).
In the second scenario, we proceed to investigate how NNN tunnelings of $t_{a}$ and $t_{b}$ modify the soliton’s transport predicted by Fig. 5. We first examine the case $t_{a}=0.4$, $t_{b}=-0.1$ and $g=1$. As depicted in Figs. 6(a), the gapless edge states persist due to the preservation of inversion symmetry, which corresponds to the stable soliton oscillations shown in Figs. 6(b). Similar to the case with linear eigenvalues, the pumping value here is 2, and there is perfect adiabatic following. When $t_{a}=-t_{b}=0.5$, the ground-state soliton exhibits a self-intersecting structure, as illustrated in Figs. 6(c). This self-intersecting band structure leads to the breakdown of the adiabatic path, as shown in Figs. 6(d). Moreover the results in Figs. 6(e)demonstrate that NNN parameters $\{t_{a},t_{b}\}$ actively reconfigure the nonlinear system’s phase diagram—a capability absent in linear eigenvalue systems. Specifically, the NNN parameters of $t_{a}=-t_{b}=0.5$ (see the green line Figs. 6(e)) induces a fractional topological phase transition distinct from linear mechanisms in Fig. 2(b).
Figure 7: Effects of parameter $t_{a}$ and $t_{b}$ on the auxiliary $\lambda$-spectrum defined in Eq. (14) for the nonlinear extended RM Hamiltonian (1).The parameters are fixed at $J=1$, $\delta=0.5$, $\Delta=1$, $g=7.5$, $T=2000\pi$, and $\omega=10^{-3}$. Other parameters are given as (a-b): $t_{a}=0.4$, $t_{b}=-0.1$; (c-d): $t_{a}=0.5$, $t_{b}=-0.5$. (e): Effects of parameter $t_{a}$ and $t_{b}$ on the anticipated position of the nonlinear excitation of soliton as a function of time over one period.
In the third scenario, we further investigate the influence of the NNN hopping terms $t_{a}$ and $t_{b}$ on the nonlinear eigenvalue system by choosing $g=7.5$. It is noted that the nonlinearity strength $g$ at this point is higher than that set for the linear eigenvalues, due to the fact that the introduction of nonlinear eigenvalues causes the critical point of the nonlinearity strength $g$, influenced by $t_{a}$ and $t_{b}$ in the phase diagram, to increase.When $t_{a}=-t_{b}=0.5$, the energy spectrum exhibits a self-intersecting structure of the ground-state soliton, as shown in Figs. 7(c). Correspondingly, in the phase diagram, adiabatic breakdown is observed, as depicted in Fig. 7(d). In contrast, when $t_{a}=0.4$, $t_{b}=-0.1$, similar to the case with linear eigenvalues, the self-intersecting structure of the ground-state soliton disappears, and the pumping effect ceases, as shown in Figs. 7(a-b).
The comparative analysis in Figs. 7(e) reveals the parametric dependence of soliton dynamics: akin to the case with linear eigenvalues, adiabatic breakdown still occurs when $t_{a}=-t_{b}$; whereas in other cases, the pumping effect is halted. It is noteworthy that, unlike the linear eigenvalue case, when $t_{a}=t_{b}=0$, the phase diagram shows the cessation of the pumping effect, whereas in the linear eigenvalue scenario, adiabatic breakdown is observed. However, it should be noted that the case where $t_{a}=t_{b}=0$ corresponds to the original nonlinear RM model, not the nonlinear extended RM model, and thus this phenomenon can be considered negligible. The above analysis indicates that the NNN terms do indeed selectively modify the system’s phase diagram, especially when the parameter $t_{a}=-t_{b}$.
This section presents the energy spectra and phase diagrams for the nonlinear extended RM model under linear and nonlinear eigenvalue conditions. Comparative analysis shows that the nonlinearity strength $g$ controls phase diagram evolution: increasing $g$ drives transitions from perfect adiabatic transport through breakdown to complete suppression of pumping. The NNN parameters $t_{a}$ and $t_{b}$ modulate these phase transitions. Crucially, when $t_{a}=-t_{b}$, the phase diagram exhibits characteristic restructuring that manifests as ground-state soliton reconfiguration in the energy spectrum. We further observe fractional topological phases emerging at specific combinations of $g$, $t_{a}$, and $t_{b}$. These phases display transport properties sensitive to eigenvalue variations, reflecting changes in topological invariants.
IV Nonlinear bulk-edge correspondence
In the preceding Sec. III, we studied the nonlinear extended RM model under linear and nonlinear eigenvalue conditions, obtaining its $\lambda$ spectrum and ground-state pumping diagram. Furthermore, we determined the impact of NNN parameters $t_{a}$ and $t_{b}$ on the system’s phase and energy spectra. Section IV extends this investigation by analyzing how $t_{a}$ and $t_{b}$ influence anomalous nonlinear eigenvalue solutions in the extended RM Hamiltonian (1).
Figure 8: Band structure $\omega$-$t$ of the auxiliary eigenvalue by solving Eq. (14) at $\lambda=0$. The parameters are fixed at $g=2$, $J=1$, $\delta=0.5$, $\Delta=1$, $T=2000\pi$, and $\omega=10^{-3}$. Parameters $t_{a}$ and $t_{b}$ are given as (a): $t_{a}=0$, $t_{b}=0$; (b): $t_{a}=0.4$, $t_{b}=0.4$; (c): $t_{a}=0.4$, $t_{b}=-0.1$; (d): $t_{a}=0.4$, $t_{b}=-0.4$.
First, we analyze how the NNN hopping parameters $t_{a}$ and $t_{b}$ affect nonlinearity in the anomalous eigenvalue problem for the extended Rice-Mele Hamiltonian (Eq. 1). We specifically solve Eq. (14) at $\lambda=0$ to examine the relationship $t$$\omega$. This is achieved by computing $t$$\omega$ diagrams for varying $t_{a}$ and $t_{b}$ at fixed $\lambda=0$. Our analysis shows two key features: (i) Soliton states emerge irrespective of the specific values of $t_{a}$ and $t_{b}$, and (ii) increasing the interaction strength $g$ lowers the ground-state soliton energy. To clearly resolve two distinct soliton states, we focus our analysis on $g=2$.
Next, we plot the $\omega$-$t$ diagrams for different $t_{a}$ and $t_{b}$ values at $\lambda=0$ and $g=2$ (Fig. 8). In the original model ($t_{a}=t_{b}=0$; Fig. 8(a)), edge states display a sharp turning point. Setting $t_{a}=t_{b}=0.4$ (Fig. 8(b)) broadens the bulk bands and introduces a new bulk band between them. For $t_{a}=0.4$, $t_{b}=-0.1$ (Fig. 8(c)), bulk band widths decrease and the intermediate band diminishes. Here, the central soliton state connects to this intermediate band, becoming more distinct as bands narrow. The upper edge state retains its sharp turn, exhibiting fracture near $\omega=0$ in the latter half-period, while the lower state curves. When $t_{a}=-t_{b}=0.4$ (Fig. 8(d)), band edges narrow at the upper bulk band while the lower band widens centrally, with complete disappearance of the intermediate band. This fully reveals the soliton band. Both edge states maintain their characteristic forms (sharp-angled upper, curved lower), with the fractured upper edge subsequently reconnecting.
At last, the NNN parameters $t_{a}$ and $t_{b}$ modify the anomalous eigenvalue solutions of the extended RM model. The most significant modifications occur in the $\omega$-$t$ diagram when $t_{a}=-t_{b}$. Specifically: (i) Nonzero $t_{a}$, $t_{b}$ values generate a new bulk band between existing bands; (ii) When $t_{a}=-t_{b}$, this emergent band vanishes entirely.
V Conclusion and outlook
This study primarily investigates the behavior of the nonlinear extended RM model under both linear and nonlinear eigenvalue conditions. The research explicitly demonstrates the energy spectra and phase diagrams of the system for these two distinct eigenvalue scenarios, emphasizing that the parameters $t_{a}$ and $t_{b}$ of the NNN terms can influence the system’s soliton behavior and pumping transport. Notably, we find that assigning certain specific values to the parameters $t_{a}$ and $t_{b}$ induces the emergence of fractional topological phases in the system. Moreover, when $t_{a}=-t_{b}$, the range of nonlinearity strength for adiabatic breakdown is expanded, meaning that the value of nonlinearity strength at which adiabatic breakdown begins is reduced, while the value at which it ceases is increased. Moreover, we have uncovered a fascinating phenomenon where the nonlinear parameter g can modulate the system’s topological invariants.
The study also establishes the bulk-edge correspondence in the anomalous nonlinear extended RM model. We find that the parameters $t_{a}$ and $t_{b}$ of the NNN terms also modify the system’s $\omega-t$ diagram. Specifically, when $t_{a}=-t_{b}$, this diagram exhibits a striking modification: the bulk bands emerging from the NNN terms vanish entirely. This work provides key insights into the extended RM model’s behavior under nonlinearity, revealing new phenomena at the interplay between nonlinear interactions and NNN couplings. These results have important implications for designing topological insulators and controlling edge states in nonlinear quantum systems, motivating further studies on nonlinear topological phenomena.
Acknowledgements – We thank Ying Hu, Yapeng Zhang, Xunzhen Cao, Shujie Cheng, and Biao Wu for stimulating discussions and useful help. We sincerely acknowledge Yucan Yan for his dedicated efforts during the critical inception phase of this project. This work was supported by the Zhejiang Provincial Natural Science Foundation (Grant No. LZ25A040004) and the National Natural Science Foundation of China (Grants No. 12074344).